OpenStudy (anonymous):

what is the domain of this function : 4/x+8

OpenStudy (anonymous):

\[1\div x+2\]

OpenStudy (anonymous):

dude the domain???

OpenStudy (anonymous):

is this = 4 / (x+8)?

OpenStudy (anonymous):

ya ....i need the domain of that doh

OpenStudy (anonymous):

thank you for the help

OpenStudy (anonymous):

if so the domain is all real numbers except -8

OpenStudy (anonymous):

ok koo could you explain how...if you can

OpenStudy (anonymous):

Sorry my English isn't that good, though you needed to simplify it.

OpenStudy (anonymous):

x=-8 give the function an indeterminate value

OpenStudy (anonymous):

4 / (-8+8) = 4 /0 which is indeterminate

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

4/0 cannot be calculated . No matter how many times you add 0's the answer is still 0 it will never reach 4

OpenStudy (anonymous):

if i could give you more medals i would

OpenStudy (anonymous):

thats ok